# Unified Evolution of Electromagnetic Sources in Homogeneous Fields

### v1

@@ IV.1 Explicit Form and Non-Relativistic Limit
$\frac{\mathrm{d(\gamma\pmb{\mu})}}{\mathrm{d}\tau}$ $=\frac{eg}{2m}\gamma(\pmb{d}\times\mathbf{E}+\pmb{\mu}\times\mathbf{B})-\frac{e(g-2)}{2mc^{2}}\gamma^{3}\Big{(}(\mathbf{v}\cdot\mathbf{E})(\mathbf{v}\times\pmb{d})+\mathbf{v}\times(\pmb{\mu}\times\mathbf{E})-(\mathbf{v}\cdot\pmb{\mu})(\mathbf{v}\times\mathbf{B}))\Big{)}\ ,$ (35)
@@
which can be reformulated to (see the Appendix for details)
@@ IV.2 Spin-Pseudovector Representation
from which the connection $s^{0}=\frac{1}{c}\mathbf{v}\cdot\mathbf{s}$ arises. Applying the effective Lorentz-like transformation on the spin-pseudovector and reconstructing the proper time derivative using the same argument as before yields the evolution equation for $s^{\mu}$ (see the Appendix for details)
@@
The above equation matches the one in literature on the evolution of the magnetic moment in terms of a four-pseudovector (see, e.g., [13, 10, 36]). The two descriptions of classical spin – the tensor $\Sigma^{\mu\nu}$ and vector $s^{\mu}$ ones – are equivalent and a dual transformation exists between them [37, Chapter 10.4], which reads as
@@
Equations (33) and (41) are compatible with the dual transformation in the above equation, and one can use it to switch between the evolution equations in the two representations. The corresponding calculation is sketched in the Appendix.
@@ Acknowledgments
I.I. acknowledges support by the Karlsruhe School of Optics & Photonics (KSOP).
@@ B.1 Infinitesimal Lorentz Transformation
$J_{x}=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&-1\\ 0&0&1&0\\ \end{pmatrix}\ ,\ \ J_{y}=\begin{pmatrix}0&0&0&0\\ 0&0&0&1\\ 0&0&0&0\\ 0&-1&0&0\\ \end{pmatrix}\ ,$ (47)
$J_{z}=\begin{pmatrix}0&0&0&0\\ 0&0&-1&0\\ 0&1&0&0\\ 0&0&0&0\\ \end{pmatrix}\ ,\ \ K_{x}=\begin{pmatrix}0&1&0&0\\ 1&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ \end{pmatrix}\ ,$
$K_{y}=\begin{pmatrix}0&0&1&0\\ 0&0&0&0\\ 1&0&0&0\\ 0&0&0&0\\ \end{pmatrix}\ ,\ \ K_{z}=\begin{pmatrix}0&0&0&1\\ 0&0&0&0\\ 0&0&0&0\\ 1&0&0&0\\ \end{pmatrix}\ .$
@@
$L^{\mu}_{\ \nu}=\delta^{\mu}_{\nu}+\begin{pmatrix}0&\zeta_{1}&\zeta_{2}&\zeta_{3}\\ \zeta_{1}&0&-\varphi_{3}&\varphi_{2}\\ \zeta_{2}&\varphi_{3}&0&-\varphi_{1}\\ \zeta_{3}&-\varphi_{2}&\varphi_{1}&0\\ \end{pmatrix}^{\mu}_{\ \ \nu}$
@@ B.3 Metric Tensor Invariance
=$\stackrel{{\scriptstyle!}}{{=}}\eta^{\mu\nu}\mathrm{d}x^{\prime}_{\mu}\mathrm{d}x^{\prime}_{\nu}$
@@
$=\big{(}\delta^{\mu}_{\beta}+\Delta\tau\alpha F^{\mu}_{\ \beta}\big{)}\big{(}\delta^{\nu}_{\sigma}-\Delta\tau\alpha F_{\sigma}^{\ \nu}\big{)}\eta^{\beta\sigma}$
$=\big{(}\delta^{\mu}_{\beta}+\Delta\tau\alpha F^{\mu}_{\ \beta}\big{)}\big{(}\eta^{\beta\nu}-\Delta\tau\alpha F^{\beta\nu}\big{)}$
$=\eta^{\mu\nu}-\Delta\tau\alpha F^{\mu\nu}+\Delta\tau\alpha F^{\mu\nu}$
@@ C.2 Effective Lorentz-like Transformation Acting on the Moment Tensor
Replacing $F^{\mu\nu}$ by ${F^{\prime}}^{\mu\nu}$ in the commutator equation for the electromagnetic moment tensor, Eq. (23) yields
$\frac{\mathrm{d}\Sigma^{\mu\nu}}{\mathrm{d}\tau}$ $=\frac{eg}{2m}\big{(}{F^{\prime}}^{\mu}_{\ \alpha}\Sigma^{\alpha\nu}-{F^{\prime}}^{\nu}_{\ \alpha}\Sigma^{\alpha\mu}\big{)}$ (55)
$=\frac{eg}{2m}\big{(}F^{\mu}_{\ \alpha}\Sigma^{\alpha\nu}-F^{\nu}_{\ \alpha}\Sigma^{\alpha\mu}\big{)}+\frac{eg}{2m}u_{\alpha}\big{(}\Sigma^{\alpha\nu}K^{\mu}-\Sigma^{\alpha\mu}K^{\nu}\big{)}-\frac{eg}{2m}K_{\alpha}\big{(}u^{\mu}\Sigma^{\alpha\nu}-u^{\nu}\Sigma^{\alpha\mu}\big{)}$
$=\frac{eg}{2m}\big{(}F^{\mu}_{\ \alpha}\Sigma^{\alpha\nu}-F^{\nu}_{\ \alpha}\Sigma^{\alpha\mu}\big{)}-\frac{eg}{2m}K_{\alpha}\big{(}u^{\mu}\Sigma^{\alpha\nu}-u^{\nu}\Sigma^{\alpha\mu}\big{)}\ ,$
@@
$\ \ \ \,-\frac{eg}{2m}K_{\alpha}\big{(}u^{\mu}\Sigma^{\alpha\nu}u_{\nu}-u^{\nu}u_{\nu}\Sigma^{\alpha\mu}\big{)}$
$=-\frac{eg}{2m}F^{\nu}_{\ \alpha}u_{\nu}\Sigma^{\alpha\mu}+\frac{egc^{2}}{2m}K_{\alpha}\Sigma^{\alpha\mu}\ .$
@@
where we have substituted the Lorentz force, Eq. (13), in the first line and used $\Sigma^{\mu\nu}=-\Sigma^{\nu\mu}$ and then switched the dummy indices $\alpha$ and $\nu$ in the second one. Equations 58, 57 and 56 can then be combined into
$\frac{e}{m}\Sigma^{\alpha\mu}\bigg{(}-\frac{g}{2}F^{\nu}_{\ \alpha}u_{\nu}+\frac{gc^{2}}{2}K_{\alpha}+F^{\nu}_{\ \alpha}u_{\nu}\bigg{)}=0\ ,$ (59)
by using that $F_{\alpha\nu}=-F_{\nu\alpha}$ and $F_{\nu\alpha}u^{\nu}=F^{\nu}_{\ \alpha}u_{\nu}$ to reformulate the right-hand side of Eq. (58). Requiring that the whole term in parenthesis in the above equation vanishes lets us determine the vector $K_{\alpha}$ to be
$K_{\alpha}=\frac{1}{gc^{2}}(g-2)F^{\nu}_{\ \alpha}u_{\nu}\ ,$ (60)
which corresponds to Eq. (32). Substituting the vector $K_{\alpha}$ in Eq. (55) and rewriting $(u^{\mu}\Sigma^{\alpha\nu}-u^{\nu}\Sigma^{\alpha\mu})F^{\nu}_{\ \alpha}u_{\nu}=(u^{\mu}\Sigma^{\nu\alpha}-u^{\nu}\Sigma^{\mu\alpha})F_{\alpha\nu}u^{\nu}$ yields Eq. (33).
@@ C.3 Effective Lorentz-like Transformation Acting on the Electric Sources
${u^{\prime}}^{\mu}$ $={\Lambda^{\prime}}^{\mu}_{\ \nu}u^{\nu}$ (61)
$=\bigg{(}\delta^{\mu}_{\nu}+\Delta\tau\frac{eg}{2m}{F^{\prime}}^{\mu}_{\ \nu}\bigg{)}u^{\nu}$
$=\bigg{(}\delta^{\mu}_{\nu}+\Delta\tau\frac{eg}{2m}\big{(}F^{\mu}_{\ \nu}+K^{\mu}u_{\nu}-K_{\nu}u^{\mu}\big{)}\bigg{)}u^{\nu}$
@@
$=u^{\mu}+\Delta\tau\frac{e}{m}{F}^{\mu}_{\ \alpha}u^{\alpha}\ ,$
where we have applied the transformation on the four-velocity, which is written first in terms of ${F^{\prime}}^{\mu\nu}$, then in terms of $K^{\mu}$ and finally in its most explicit but still covariant form in the first four steps of the above equation. Opening the parenthesis in the next line leads to the vanishing of the last term since $F_{\alpha\nu}u^{\alpha}u^{\nu}=0$. The property is easily obtained when one considers that on the one hand $F_{\alpha\nu}u^{\alpha}u^{\nu}=F_{\nu\alpha}u^{\nu}u^{\alpha}$ when simply exchanging the dummy indices $\alpha$ and $\nu$ and on the other hand $F_{\alpha\nu}u^{\alpha}u^{\nu}=-F_{\nu\alpha}u^{\alpha}u^{\nu}$ upon utilizing the total asymmetry of the electromagnetic tensor. Equating both results yields $F_{\nu\alpha}u^{\nu}u^{\alpha}=-F_{\nu\alpha}u^{\alpha}u^{\nu}$, which means that the expression can only be identically zero. In the sixth line we have also used that $u_{\mu}u^{\mu}=c^{2}$ in the third term as well as the reformulation $F^{\alpha\mu}u_{\alpha}=-F^{\mu\alpha}u_{\alpha}=-F^{\mu}_{\ \alpha}u^{\alpha}$. Finally, we see that the second term and the part of the third one, which is proportional to $g$, are identical up to a sign and hence cancel. We are left with the result in the last line, which yields the Lorentz force directly upon the use of Eq. (18) and reconstruction of the proper time derivative.
@@ C.5 Effective Lorentz-like Transformation Acting on the Spin-Pseudovector
Applying the effective Lorentz-like transformation on the spin-pseudovector according to the transformation rule for a four-vector and writing down the transformation in terms of ${F^{\prime}}$ yields
${s^{\prime}}^{\mu}$ $={\Lambda^{\prime}}^{\mu}_{\ \nu}s^{\nu}$ (68)
$=\bigg{(}\delta^{\mu}_{\nu}+\Delta\tau\frac{eg}{2m}{F^{\prime}}^{\mu}_{\ \nu}\bigg{)}s^{\nu}$
$=\bigg{(}\delta^{\mu}_{\nu}+\Delta\tau\frac{eg}{2m}\big{(}F^{\mu}_{\ \nu}+K^{\mu}u_{\nu}-K_{\nu}u^{\mu}\big{)}\bigg{)}s^{\nu}$
@@
Here, the third line $F^{\prime}$ has been expanded in terms of $K$, whose form has been written explicitly in the next step. Finally, the condition in Eq. (40) has been used to drop the third term in the parentheses, resulting in the final form. Reconstructing the proper time derivative by taking ${s^{\prime}}^{\mu}\rightarrow s^{\mu}(\tau+\Delta\tau)$ then leads to Eq. (41)
@@ Appendix D Dual Transformation Between the Moment Tensor and Pseudovector
The following steps were made in the equation above. In the second line we have used that $\epsilon_{\alpha\beta\mu\nu}=\epsilon_{\mu\nu\alpha\beta}$ and substituted Eq. (69b), and have then applied the emerging Kronecker deltas to the term $(s_{\delta}u_{\sigma})$ in the next one. In the fourth step, the product rule has been performed for the proper time derivative. In the following one, we have used the fact that $s_{\beta}u^{\beta}=0$ and $\frac{\mathrm{d}u_{\beta}}{\mathrm{d}\tau}u^{\beta}=0$, i.e., Eqs. (40) and (4), to remove the second and fourth term in the parenthesis as well as the norm of the four-velocity in the third term, $u_{\beta}u^{\beta}=c^{2}$. In the next to last line we have reformulated the second term by taking the proper time derivative of Eq. (40) and getting $\frac{\mathrm{d}s_{\beta}}{\mathrm{d}\tau}u^{\beta}=-s_{\beta}\frac{\mathrm{d}u^{\beta}}{\mathrm{d}\tau}$. In the last line we have substituted $\frac{\mathrm{d}u^{\beta}}{\mathrm{d}\tau}$ using the Lorentz force, Eq. (13) and reformulating the term as $F^{\beta}_{\ \nu}u^{\nu}s_{\beta}=-F_{\nu\beta}u^{\nu}s^{\beta}$.
We now perform an analogous set of operations to the first term on the right-hand side of Eq. (33)
@@
the result of which corresponds to Eq. (41) for the contravariant vector $s_{\mu}$. To get to the final expression, the dummy indices in the last term on the right-hand side of the first line have been exchanged, which makes obvious the fact that the last two terms are identical up to a prefactor, which, when put together, gives the desired $(g-2)$-term in the result. The conversion of the equation of motion for the magnetic sources from the $s^{\mu}$ to the $\Sigma^{\mu\nu}$ representation works in an equivalent way.

### v2

@@ IV.1 Explicit Form and Non-Relativistic Limit
$\frac{\mathrm{d(\gamma\pmb{\mu})}}{\mathrm{d}\tau}$ $=\frac{eg}{2m}\gamma(\pmb{d}\times\mathbf{E}+\pmb{\mu}\times\mathbf{B})-\frac{e(g-2)}{2mc^{2}}\gamma^{3}\Big{(}(\mathbf{v}\cdot\mathbf{E})(\mathbf{v}\times\pmb{d})+\mathbf{v}\times(\pmb{\mu}\times\mathbf{E})-(\mathbf{v}\cdot\pmb{\mu})(\mathbf{v}\times\mathbf{B})\Big{)}\ ,$ (38)
@@
which can be reformulated to (see Appendix C.4 for details)
@@ IV.2 Spin-Pseudovector Representation
from which the connection $s^{0}=\frac{1}{c}\mathbf{v}\cdot\mathbf{s}$ arises. Applying the effective Lorentz-like transformation on the spin-pseudovector and reconstructing the proper time derivative using the same argument as before yields the evolution equation for $s^{\mu}$ (see Appendix C.5 for details)
@@
The above equation matches the one in literature on the evolution of the magnetic moment in terms of a four-pseudovector (see, e.g., [13, 10, 37]). The two descriptions of classical spin – the tensor $\Sigma^{\mu\nu}$ and vector $s^{\mu}$ ones – are equivalent and a dual transformation exists between them [38, Chapter 10.4], which reads as
@@
Equations (33) and (44) are compatible with the dual transformation in the above equation, and one can use it to switch between the evolution equations in the two representations. The corresponding calculation is sketched in Appendix D.
@@ Acknowledgments
We acknowledge support by the KIT Publication Fund of the Karlsruhe Institute of Technology. I.I. acknowledges support by the Karlsruhe School of Optics & Photonics (KSOP) and the Ministry of Science, Research and Arts of Baden-Württemberg as part of the sustainability financing of the projects of the Excellence Initiative II.
@@ B.1 Infinitesimal Lorentz Transformation
$J_{x}=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&-1\\ 0&0&1&0\\ \end{pmatrix}\ ,\ \ J_{y}=\begin{pmatrix}0&0&0&0\\ 0&0&0&1\\ 0&0&0&0\\ 0&-1&0&0\\ \end{pmatrix}\$, (50)
$J_{z}=\begin{pmatrix}0&0&0&0\\ 0&0&-1&0\\ 0&1&0&0\\ 0&0&0&0\\ \end{pmatrix}\ ,\ \ K_{x}=\begin{pmatrix}0&1&0&0\\ 1&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ \end{pmatrix}\$,
$K_{y}=\begin{pmatrix}0&0&1&0\\ 0&0&0&0\\ 1&0&0&0\\ 0&0&0&0\\ \end{pmatrix}\ ,\ \ K_{z}=\begin{pmatrix}0&0&0&1\\ 0&0&0&0\\ 0&0&0&0\\ 1&0&0&0\\ \end{pmatrix}\$.
@@
$L^{\mu}_{\ \nu}=\delta^{\mu}_{\nu}+\begin{pmatrix}0&\zeta_{1}&\zeta_{2}&\zeta_{3}\\ \zeta_{1}&0&-\varphi_{3}&\varphi_{2}\\ \zeta_{2}&\varphi_{3}&0&-\varphi_{1}\\ \zeta_{3}&-\varphi_{2}&\varphi_{1}&0\\ \end{pmatrix}$
@@ B.3 Metric Tensor Invariance
$\stackrel{{\scriptstyle!}}{{=}}\eta^{\mu\nu}\mathrm{d}x^{\prime}_{\mu}\mathrm{d}x^{\prime}_{\nu}$
@@
$=\big{(}\delta^{\mu}_{\beta}+\Delta\tau\frac{e}{m}F^{\mu}_{\ \beta}\big{)}\big{(}\delta^{\nu}_{\sigma}-\Delta\tau\frac{e}{m}F_{\sigma}^{\ \nu}\big{)}\eta^{\beta\sigma}$
$=\big{(}\delta^{\mu}_{\beta}+\Delta\tau\frac{e}{m}F^{\mu}_{\ \beta}\big{)}\big{(}\eta^{\beta\nu}-\Delta\tau\frac{e}{m}F^{\beta\nu}\big{)}$
$=\eta^{\mu\nu}-\Delta\tau\frac{e}{m}F^{\mu\nu}+\Delta\tau\frac{e}{m}F^{\mu\nu}$
@@ C.2 Effective Lorentz-like Transformation Acting on the Moment Tensor
Replacing $F^{\mu\nu}$ by ${\tilde{F}}^{\mu\nu}$ in the commutator equation for the electromagnetic moment tensor, Eq. (23) yields
$\frac{\mathrm{d}\Sigma^{\mu\nu}}{\mathrm{d}\tau}$ $=\frac{eg}{2m}\big{(}{\tilde{F}}^{\mu}_{\ \alpha}\Sigma^{\alpha\nu}-{\tilde{F}}^{\nu}_{\ \alpha}\Sigma^{\alpha\mu}\big{)}$ (58)
$=\frac{eg}{2m}\big{(}F^{\mu}_{\ \alpha}\Sigma^{\alpha\nu}-F^{\nu}_{\ \alpha}\Sigma^{\alpha\mu}\big{)}+\frac{eg}{2m}u_{\alpha}\big{(}\Sigma^{\alpha\nu}\Xi^{\mu}-\Sigma^{\alpha\mu}\Xi^{\nu}\big{)}-\frac{eg}{2m}\Xi_{\alpha}\big{(}u^{\mu}\Sigma^{\alpha\nu}-u^{\nu}\Sigma^{\alpha\mu}\big{)}$
$=\frac{eg}{2m}\big{(}F^{\mu}_{\ \alpha}\Sigma^{\alpha\nu}-F^{\nu}_{\ \alpha}\Sigma^{\alpha\mu}\big{)}-\frac{eg}{2m}\Xi_{\alpha}\big{(}u^{\mu}\Sigma^{\alpha\nu}-u^{\nu}\Sigma^{\alpha\mu}\big{)}\ ,$
@@
$\ \ \ \,-\frac{eg}{2m}\Xi_{\alpha}\big{(}u^{\mu}\Sigma^{\alpha\nu}u_{\nu}-u^{\nu}u_{\nu}\Sigma^{\alpha\mu}\big{)}$
$=-\frac{eg}{2m}F^{\nu}_{\ \alpha}u_{\nu}\Sigma^{\alpha\mu}+\frac{egc^{2}}{2m}\Xi_{\alpha}\Sigma^{\alpha\mu}\ .$
@@
where we have substituted the Lorentz force, Eq. (13), in the first line and used $\Sigma^{\mu\nu}=-\Sigma^{\nu\mu}$ and then switched the dummy indices $\alpha$ and $\nu$ in the second one. Equations 61, 60 and 59 can then be combined into
$\frac{e}{m}\Sigma^{\alpha\mu}\bigg{(}-\frac{g}{2}F^{\nu}_{\ \alpha}u_{\nu}+\frac{gc^{2}}{2}\Xi_{\alpha}+F^{\nu}_{\ \alpha}u_{\nu}\bigg{)}=0\ ,$ (62)
by using that $F_{\alpha\nu}=-F_{\nu\alpha}$ and $F_{\nu\alpha}u^{\nu}=F^{\nu}_{\ \alpha}u_{\nu}$ to reformulate the right-hand side of Eq. (61). Requiring that the whole term in parenthesis in the above equation vanishes lets us determine the vector $\Xi_{\alpha}$ to be
$\Xi_{\alpha}=\frac{1}{gc^{2}}(g-2)F^{\nu}_{\ \alpha}u_{\nu}\ ,$ (63)
which corresponds to Eq. (32). Substituting the vector $\Xi_{\alpha}$ in Eq. (58) and rewriting $(u^{\mu}\Sigma^{\alpha\nu}-u^{\nu}\Sigma^{\alpha\mu})F^{\nu}_{\ \alpha}u_{\nu}=(u^{\mu}\Sigma^{\nu\alpha}-u^{\nu}\Sigma^{\mu\alpha})F_{\alpha\nu}u^{\nu}$ yields Eq. (33).
@@ C.3 Effective Lorentz-like Transformation Acting on the Electric Sources
${u^{\prime}}^{\mu}$ $={\tilde{\Lambda}}^{\mu}_{\ \nu}u^{\nu}$ (64)
$=\bigg{(}\delta^{\mu}_{\nu}+\Delta\tau\frac{eg}{2m}{\tilde{F}}^{\mu}_{\ \nu}\bigg{)}u^{\nu}$
$=\bigg{(}\delta^{\mu}_{\nu}+\Delta\tau\frac{eg}{2m}\big{(}F^{\mu}_{\ \nu}+\Xi^{\mu}u_{\nu}-\Xi_{\nu}u^{\mu}\big{)}\bigg{)}u^{\nu}$
@@
$=u^{\mu}+\Delta\tau\frac{e}{m}{F}^{\mu}_{\ \nu}u^{\nu}\ ,$
where we have applied the transformation on the four-velocity, which is written first in terms of ${\tilde{F}}^{\mu\nu}$, then in terms of $\Xi^{\mu}$ and finally in its most explicit but still covariant form in the first four steps of the above equation. Opening the parenthesis in the next line leads to the vanishing of the last term since $F_{\alpha\nu}u^{\alpha}u^{\nu}=0$. The property is easily obtained when one considers that on the one hand $F_{\alpha\nu}u^{\alpha}u^{\nu}=F_{\nu\alpha}u^{\nu}u^{\alpha}$ when simply exchanging the dummy indices $\alpha$ and $\nu$ and on the other hand $F_{\alpha\nu}u^{\alpha}u^{\nu}=-F_{\nu\alpha}u^{\alpha}u^{\nu}$ upon utilizing the total asymmetry of the electromagnetic tensor. Equating both results yields $F_{\nu\alpha}u^{\nu}u^{\alpha}=-F_{\nu\alpha}u^{\alpha}u^{\nu}$, which means that the expression can only be identically zero. In the sixth line we have also used that $u_{\mu}u^{\mu}=c^{2}$ in the third term as well as the reformulation $F^{\alpha\mu}u_{\alpha}=-F^{\mu\alpha}u_{\alpha}=-F^{\mu}_{\ \alpha}u^{\alpha}$. Finally, we see that the second term and the part of the third one, which is proportional to $g$, are identical up to a sign and hence cancel. We are left with the result in the last line, which yields the Lorentz force directly upon the use of Eq. (18) and reconstruction of the proper time derivative. In short, it holds that $\frac{e}{m}{F}^{\mu}_{\ \nu}u^{\nu}=\frac{eg}{2m}{\tilde{F}}^{\mu}_{\ \nu}u^{\nu}$.
@@ C.5 Effective Lorentz-like Transformation Acting on the Spin-Pseudovector
Applying the effective Lorentz-like transformation on the spin-pseudovector according to the transformation rule for a four-vector and writing down the transformation in terms of $\tilde{F}^{\mu}_{\ \nu}$} yields
${s^{\prime}}^{\mu}$ $={\tilde{\Lambda}}^{\mu}_{\ \nu}s^{\nu}$ (71)
$=\bigg{(}\delta^{\mu}_{\nu}+\Delta\tau\frac{eg}{2m}{\tilde{F}}^{\mu}_{\ \nu}\bigg{)}s^{\nu}$
$=\bigg{(}\delta^{\mu}_{\nu}+\Delta\tau\frac{eg}{2m}\big{(}F^{\mu}_{\ \nu}+\Xi^{\mu}u_{\nu}-\Xi_{\nu}u^{\mu}\big{)}\bigg{)}s^{\nu}$
@@
Here, the third line $\tilde{F}^{\mu}_{\ \nu}$ has been expanded in terms of $\Xi^{\mu}$, whose form has been written explicitly in the next step. Finally, the condition in Eq. (43) has been used to drop the third term in the parentheses, resulting in the final form. Reconstructing the proper time derivative by taking ${s^{\prime}}^{\mu}\rightarrow s^{\mu}(\tau+\Delta\tau)$ then leads to Eq. (44)
@@ Appendix D Dual Transformation Between the Moment Tensor and Pseudovector
The following steps were made in the equation above. In the second line we have used that $\epsilon_{\alpha\beta\mu\nu}=\epsilon_{\mu\nu\alpha\beta}$ and substituted Eq. (72b), and have then applied the emerging Kronecker deltas to the term $(s_{\delta}u_{\sigma})$ in the next one. In the fourth step, the product rule has been performed for the proper time derivative. In the following one, we have used the fact that $s_{\beta}u^{\beta}=0$ and $\frac{\mathrm{d}u_{\beta}}{\mathrm{d}\tau}u^{\beta}=0$, i.e., Eqs. (43) and (4), to remove the second and fourth term in the parenthesis as well as the norm of the four-velocity in the third term, $u_{\beta}u^{\beta}=c^{2}$. In the next to last line we have reformulated the second term by taking the proper time derivative of Eq. (43) and getting $\frac{\mathrm{d}s_{\beta}}{\mathrm{d}\tau}u^{\beta}=-s_{\beta}\frac{\mathrm{d}u^{\beta}}{\mathrm{d}\tau}$. In the last line we have substituted $\frac{\mathrm{d}u^{\beta}}{\mathrm{d}\tau}$ using the Lorentz force, Eq. (13) and reformulating the term as $F^{\beta}_{\ \nu}u^{\nu}s_{\beta}=-F_{\nu\beta}u^{\nu}s^{\beta}$. We now perform an analogous set of operations to the first term on the right-hand side of Eq. (33)
@@
the result of which corresponds to Eq. (44) for the contravariant vector $s_{\mu}$. To get to the final expression, the dummy indices in the last term on the right-hand side of the first line have been exchanged, which makes obvious the fact that the last two terms are identical up to a prefactor, which, when put together, gives the desired (g2)(g-2)-term in the result. The conversion of the equation of motion for the magnetic sources from the $s^{\mu}$ to the $\Sigma^{\mu\nu}$ representation works in an equivalent way.